OpenStudy (anonymous):

Which equation illustrates the additive inverse property? A. c + d = d + c B. c × d = d × c C. -c + (-c) = 0 D. d + (-d) = 0

OpenStudy (anonymous):

@lily_980

OpenStudy (anonymous):

@TriniEmo

OpenStudy (anonymous):

it's D.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

do you know this one? @angel12310 Which equation illustrates the additive inverse property? A. -3 + 3 = 3 + (-3) B. 3 × 0 = 0 C. (3 + 0) + 1 = 3 + (0 + 1) D. -3 + 3 = 0

OpenStudy (anonymous):

it's A.

OpenStudy (anonymous):

thnx!!

OpenStudy (anonymous):

welcome